OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 9-inch perimeter

OpenStudy (anonymous):

area =[under root3/4] *(side)^2 =[under root3/4]*81 =35.07

OpenStudy (anonymous):

so like i said before sqrtanswer

OpenStudy (anonymous):

ok//=[(sqrt3)/4] *81 =81 (sqrt3)/4

OpenStudy (anonymous):

right?